OpenStudy (anonymous):

My TWO POINTS ARE---> (0,5) and (2,3) 1.Select two ordered pair points of your own. (Just pick any two points that seem relatively easy to work with.) Be sure to list these points for your instructor. Find the slope and show your work. 2.Take the points you have selected and write the equation of the line using the point-slope formula. Change the equation to slope-intercept form. Show the work. 3.Take the points you have selected and write the equation of the line using the slope-intercept formula. Show the work.

OpenStudy (anonymous):

please help me ive tried so many different ways and it never works

OpenStudy (kropot72):

Do you know the formula to find the slope from two given points?

OpenStudy (anonymous):

yeah y2-y1/x2-x1? or y-y1=m(x-x1)

OpenStudy (kropot72):

What value did you get for the slope?

OpenStudy (anonymous):

1

OpenStudy (kropot72):

Not really. \[\frac{y _{2}-y _{1}}{x _{2}-x _{1}}=\frac{3-5}{2-0}=\ ?\]

OpenStudy (anonymous):

i did that and got -2/2 if you divide that it equals 1

OpenStudy (anonymous):

or -1

OpenStudy (kropot72):

The result -1 is correct. The result 1 is what you get from 2/2. Do you understand?

OpenStudy (anonymous):

yes thank you also is #1 done?

OpenStudy (kropot72):

Yes, that finishes #1. Do you need help with #2?

OpenStudy (anonymous):

yes and three please

OpenStudy (kropot72):

The standard form of the equation of a line is y = mx + b. If we plug in the value found for m we get y = -1x + b which can be simplified to y = -x + b Do you follow so far?

OpenStudy (anonymous):

yes go on omg thanks so much

OpenStudy (anonymous):

so write that for two and move to three? i just dont get the changing from point slope to slope intercept

OpenStudy (kropot72):

The question asks you to use the point-slope formula which is \[y-y _{1}=m(x-x _{1})\] Plugging in the values for the point (0, 5) gives y - 5 = -1(x - 0) which simplifies to y - 5 = -x ........(1) The slope intercept form is y = mx + b. If you add 5 to both sides of equation (1) you will get the required solution. Can you do that?

OpenStudy (anonymous):

um...... i have no idea, y+0=-5x?

OpenStudy (kropot72):

y - 5 = -x ........(1) Adding 5 to both sides of equation (1) gives y - 5 + 5 = -x + 5 ..........(2) Can you now simplify (2)?

OpenStudy (anonymous):

i have to go in 10 minutes and yes it will be y-10=-5x?

OpenStudy (kropot72):

y - 5 + 5 = -x + 5 ..........(2) The result of simplifying the left hand side is y - 5 + 5 = y + 0 = y Therefore (2) simplifies to y = -x + 5

OpenStudy (anonymous):

so thats the answer for two?

OpenStudy (kropot72):

Yes. I think we have covered the work for #3. I have to go now also.